The table below lists initial velocities measured
Answer #1

From the given chart firstly we will find out the reciprocal values to graph a lineweaver-burk plot. As:

(a) Now we graph the plot as uploaded here:

(b) The apparent values of KM and Vmax are as can calculated:

   - 1/KM = x intercept

    so from the graph       - 1/KM = - 0.5   (when inhibitor is absent)

                                       KM   = -1/-0.5 = 2mM

                                           -1/KM = - 0.2     (inhibitor present)

                                       KM = -1/-0.2 = 5mM

        1 / Vmax = y intercept

                             1/ Vmax = 1/0.10   = 10 molecules/min

In a competetive enzyme inhibition Vmax does not change with the presence of inhibitor while KM increases as shown above.

(c) In a competetive inhibition,at a given moment, the enzyme may be bound to the inhibitor, the substrate, or neither, but it can bind both at the same time. A competetive inhibitor could bind to an allosteric site of the free enzyme to prevent the substrate binding, as long as it does not bind to the allosteric site when the substrate is bound.